# Powerformer: A Transformer with Weighted Causal Attention for Time-series Forecasting

### v1

@@ Abstract
Transformers have recently shown strong performance in time-series forecasting, but their all-to-all attention mechanism overlooks the (temporal) causal and often (temporally) local nature of data. We introduce Powerformer, a novel Transformer variant that replaces noncausal attention weights with causal weights that are reweighted according to a smooth heavy-tailed decay. This simple yet effective modification endows the model with an inductive bias favoring temporally local dependencies, while still allowing sufficient flexibility to learn the unique correlation structure of each dataset. Our empirical results demonstrate that Powerformer not only achieves state-of-the-art accuracy on public time-series benchmarks, but also that it offers improved interpretability of attention patterns. Our analyses show that the model’s locality bias is amplified during training, demonstrating an interplay between time-series data and power-law-based attention. These findings highlight the importance of domain-specific modifications to the Transformer architecture for time-series forecasting, and they establish Powerformer as a strong, efficient, and principled baseline for future research and real-world applications.
Figure 1: Illustration of Powerformer and the Weighted Causal Multihead Attention (WCMHA) architecture, as well as their effects on attention weights. Panel (a) shows the Powerformer architecture (left) and the WCMHA (right). Panels (b) and (c) show the attention weights without and with our local-causal mask, respectively. Here, $\Sigma$ corresponds to the softmax function.
@@ 1 Introduction
With the recent popularity of Transformer models [36], large language models (LLMs), and foundation models [4], the Transformer architecture gained popularity in time-series forecasting, leading to a number of Transformer based state-of-the-art models: Autoformer [42], FEDformer [48], and PatchTST [27].
The multihead attention (MHA) mechanism is at the core of these models, which is an all-to-all process that enriches each embedding via a weighted combination of other embeddings based on a similarity calculation. For natural language processing (NLP) tasks, each embedding represents a word. Here, an all-to-all embedding mechanism is natural since correlations between words can be only weakly dependent on their ordering and the number of words between them. However, for time-series data, where each embedding is a temporal measurement, the relation between embeddings is both causal and dependent on the time delay. This disparity between the attention mechanism’s all-to-all structure motivates the open question: how well-suited is the Transformer’s attention mechanism for time-series data? For example, [44] shows that linear models outperform Transformers on common time-series benchmarks. Still, Transformers have seen success in time-series forecasting.
Previous methods sought to impose causal and local implicit biases on Transformer. Before Transformer’s widespread adoption, methods such as WaveNet [34] and [12] used causal convolutions. Imparting local and causal implicit biases to MHA requires altering its architecture. Reformer [19] calculates the attention weights via a causal and locality-aware hash map. LogSparse self-attention [20] uses convolutions to encode local variations and feeds them into a causally sparse self-attention mechanism. The work most similar to ours is ETSformer [39], which develops exponential smoothing attention that replaces the similarity-based attention weights with exponential decaying temporal weights. These temporal weights are no longer dependent on a similarity metric like the key-query projection. ETSformer also employs a frequency attention scheme alongside their exponentially weighted attention.
When enforcing a locality bias, previous methods imposed temporal length scales on the dynamics and/or the functional form of the temporal correlations. However, the temporal correlations between data points vary in both functional form and time scale between datasets. Temporal convolutions, and their dilations, are examples of imposing such a time scale. Exponentially weighted attention imposes exponentially decaying temporal correlations but lacks the similarity-weighted MHA that captures varying temporal dependencies. Previous successes from adding local and causal biases [39, 19, 20, 34, 12] show their importance, but the application method is crucial.
In this work, we develop Powerformer (illustrated in Fig. 1), a Transformer-based model that uses Weighted Causal Multihead Attention (WCMHA) to learn temporal dependencies unique to each dataset. Our model is motivated by the observation that many physical systems exhibit heavy-tailed autocorrelations, e.g., the pairwise correlation strength may decay as a power law distribution, as the time delay grows [9]. Our goal is to impose this power-law temporal dependence as an inductive bias, while allowing Powerformer the flexibility to perturb it sufficiently to capture the temporal pairwise correlations unique to each dataset. To do so, we add a temporally decaying mask to the attention mechanism, specifically to the key-query overlap (referred to as the similarity score). The mask decays attention weights and pairwise dependencies to resemble a power law, but its additive nature maintains the key and query similarity score. This maintained similarity score allows the model to learn data-dependent perturbations to the power-law pairwise dependency. The mask acts as a regularizer that guides the attention weight temporal dependence towards a power law, but still maintains sufficient expressivity for MHA to learn pairwise dependencies unique to each dataset. Moreover, Powerformer’s simplicity promotes interpretability studies that provide evidence of learned heavy-tailed dependencies in the attention weights.
We build Powerformer by combining WCMHA, the commonly used encoder-only Transformer architecture, and patching [27]. Powerformer outperforms previous state-of-the-art models on typical tasks despite its simplicity compared to other attention mechanisms. This suggests that these biases and our implementation better align Transformer with natural time-series structures.
@@
  • We develop Powerformer, a Transformer-based model for time-series, with WCMHA to impose causal and local implicit biases that shape attention’s pairwise correlation to better respect the data’s unique temporal structure.
  • Powerformer is powered by various weighting schemes: power-law decays and Butterworth filters. The former resembles naturally occurring time-series, while the latter acts analogously to a step function.
  • Our empirical results demonstrate that Powerformer obtains state-of-the-art performance when compared to similar Transformer models on a range of public times-series benchmarks. Moreover, we provide interpretability studies that show how our implicit biases shape Powerformer’s learned attention representations.
  • Powerformer can reduce the attention complexity from quadratic to linear by defining a cutoff time after which the attention weighting removes their contribution.
@@ 2 Related Work
The importance of time-series data has led to decades of work, resulting in a myriad of methods: early statistical methods [17, 18, 33, 38] (such as ARIMA [5]); multilayer perceptions (MLP) [44, 21, 10, 46] (such as NHITS [7], TSMixer [8], and N-BEATS [28]); convolutional neural networks (CNNs) [15, 12, 3] (such as TimesNet [40] and SCINet [22]); recurrent neural networks (RNNs) [31, 16] (such as DeepAR [30]); Transformers [49, 13, 39, 48, 26, 41, 43, 23, 47, 19, 20, 32, 11] (such as PatchTST [27], Autoformer [42], FEDformer [48], and iTransformer [24]); and state space models (SSMs) [1, 37, 45] (such as MAMBA [14]). Of these methods, we focus on Transformers.
The success of LLMs and the introduction of powerful foundation models, like GPT-2 [29], inspired models built atop them and others looking to improve upon the Transformer architecture. For example, Time-GPT [13] uses the Transformer encoder-decoder structure, and One-Fits-All [49] is built atop a pretrained GPT-2 and fine-tunes the attention layer normalization and temporal embedding. Some models like AutoTimes [25] leverage LLM’s in-context learning capabilities and employ textual prompt engineering to feed into LLMs. Other models have adapted the attention mechanism to time-series data: some models strongly rely on, or partially include, Fourier representations when calculating attention weights [48, 39]; AutoFormer [42] leverages temporal autocorrelations as a similarity metric; Reformer [19] uses a locality-sensitive hashing mechanism to reduce the attention’s complexity; Informer [47] selects only high attention weight values, increasing prediction capacity; FlowFormer [41] uses flow conservation to achieve linear complexity attention without imposing locality implicit biases; and iTransformer [24] inverts the temporal and embedding dimensions, applying attention along the embedding.
Transformers were developed for discrete inputs, unlike the continuous valued time-series inputs. Patching along the time domain (PatchTST) shortens the context window which allows for longer lookback windows and significantly improves performance over more complicated Transformer methods [27]. This simple yet powerful improvement has made patching ubiquitous. Other works look to convert time-series into discretized tokens similar to word vectors. Chronos [2] normalizes the input sequences and matches these continuous values to tokens made from discretizing a sufficiently large domain. TOTEM [32] similarly discretizes the input time series, but instead VQVAE [35] encodes the input and matches it with the most similar token.
@@ 3 Method
We propose Weighted Causal Multihead Attention (WCMHA) as a foundation for Powerformer, a Transformer-based architecture designed for time-series forecasting. In this section, we review standard Multihead Attention (MHA) and demonstrate how we introduce locality and causality into it. We then present various weighting functional forms and insights on how WCMHA modifies attention distributions. We then introduce the Powerformer architecture that builds upon these principles. Figure 1 provides a high-level illustration of WCMHA and its effect on attention weights.
3.1 Weighted Causal Multihead Attention
Transformers typically leverage MHA to compute new embeddings from weighted sums of input embeddings. Let $\mathbf{X}\in\mathbb{R}^{T\times d}$ be a sequence of $T$ input vectors, each of dimension $d$. For each attention head $h$, MHA first projects $\mathbf{X}$ into query, key, and value matrices, respectively denoted by $\mathbf{Q}_{h}$, $\mathbf{K}_{h}$, and $\mathbf{V}_{h}$. Concretely,
$\mathbf{Q}_{h}=\mathbf{X}\,\mathbf{W}_{h}^{(\text{Q})},\quad\mathbf{K}_{h}=\mathbf{X}\,\mathbf{W}_{h}^{(\text{K})},\quad\mathbf{V}_{h}=\mathbf{X}\,\mathbf{W}_{h}^{(\text{V})},$ (1)
where each projection matrix $\mathbf{W}_{h}^{(\cdot)}\in\mathbb{R}^{d\times d_{k}}$ is learnable and $d_{k}$ is the dimensionality of each head. The parameters $\mathbf{Q}_{h}$, $\mathbf{K}_{h}$, and $\mathbf{V}_{h}$ thus each have shape $T\times d_{k}$.
The unnormalized attention similarity scores $\mathbf{S}_{h}$ measure how each query interacts with all keys:
$\mathbf{S}_{h}=\frac{\mathbf{K}_{h}\,\mathbf{Q}_{h}^{\top}}{\sqrt{d_{k}}}.$ (2)
The factor of $1/\sqrt{d_{k}}$ helps stabilize gradients by normalizing the dot products according to the head dimension. We then apply a softmax function to obtain the normalized attention weights:
$\mathbf{C}_{h}=\text{Softmax}\bigl{(}\mathbf{S}_{h}\bigr{)},$ (3)
where each row of $\mathbf{C}_{h}$ sums to 1. Finally, each head’s output $\tilde{\mathbf{X}}_{h}$ is computed as a weighted sum of the values:
$\tilde{\mathbf{X}}_{h}=\mathbf{C}_{h}\,\mathbf{V}_{h}.$ (4)
To form the final MHA output, we concatenate the outputs from all $H$ heads and apply a linear projection:
@@ 3.1 Weighted Causal Multihead Attention
where $\mathbf{W}^{(\text{A})}\in\mathbb{R}^{(H\,d_{k})\times d}$ is another learnable matrix. The result $\tilde{\mathbf{X}}$ has the same dimensionality as the input $\mathbf{X}$ (i.e., $T\times d$), allowing it to be passed into subsequent Transformer layers or decoded for downstream tasks.
3.1.1 Enforcing Causality
To ensure that information at time $t$ does not leak from future time steps $t^{\prime}>t$, we apply a causal mask $\mathbf{M}^{(\text{C})}$:
$\mathbf{M}^{(\text{C})}_{i,j}=\begin{cases}-\infty&j>i,\\ 0&j\leq i.\end{cases}$ (6)
This mask is added to the attention score $\mathbf{S}_{h}$, forcing any attention weight for $j>i$ to be zero after the softmax:
@@ 3.1.1 Enforcing Causality
Hence, $\mathbf{C}_{t,t^{\prime}}=0$ for all $t^{\prime}>t$, ensuring causal structure for time-series data.
3.1.2 Incorporating Locality via Decaying Masks
Time-series often exhibit stronger correlations between nearby time steps compared to distant ones [9]. To bias the model toward local dependencies, we introduce a second mask $\mathbf{M}^{(\text{D})}$ that decays attention weights as the time gap $|\Delta t|$ increases:
$\mathbf{M}^{(\text{D})}_{i,j}=\begin{cases}0&j>i,\\ f(t_{i}-t_{j})&j\leq i,\end{cases}$ (7)
where $f(\Delta t)\leq 0$ is a non-positive function of the time difference. Adding $\mathbf{M}^{(\text{D})}$ to $\mathbf{S}_{h}^{(\text{C})}$ yields
$\mathbf{S}_{h}^{(\text{C,D})}=\mathbf{S}_{h}+\mathbf{M}^{(\text{C})}+\mathbf{M}^{(\text{D})}.$ (8)
We thus obtain the Weighted Causal Attention Weights
$\mathbf{C}_{h}^{(\text{C,D})}=\text{Softmax}\bigl{(}\mathbf{S}_{h}^{(\text{C,D})}\bigr{)},$ (9)
where,
$C_{i,j}^{(\text{C,D})}\propto\begin{cases}0&j>i,\\ \exp\bigl{[}f(t_{i}-t_{j})\bigr{]}&j\leq i.\end{cases}$ (10)

### v2

# Recency Biased Causal Attention for Time-series Forecasting

@@ Abstract
Recency bias is a useful inductive prior for sequential modeling: it emphasizes nearby observations and can still allow longer-range dependencies. Standard Transformer attention lacks this property, relying on all-to-all interactions that overlook the causal and often local structure of temporal data. We propose a simple mechanism to introduce recency bias by reweighting attention scores with a smooth heavy-tailed decay. This adjustment strengthens local temporal dependencies without sacrificing the flexibility to capture broader and data-specific correlations. We show that recency-biased attention consistently improves sequential modeling, aligning Transformer more closely with the read–ignore–write operations of RNNs. Finally, we demonstrate that our approach achieves competitive and often superior performance on challenging time-series forecasting benchmarks.
@@ 1 INTRODUCTION
Recency bias is a useful inductive prior for sequential modeling: it emphasizes nearby information, while still allowing longer-range dependencies. Recurrent neural networks (RNNs) naturally embody this bias through their local and causal structure, which enables them to excel at tasks that require selective reading, ignoring, or writing of information. For example, in the flip-flop task (Liu et al., 2023), RNNs easily learn to store and discard information as needed, whereas Transformers often fail despite their greater expressivity. This limitation highlights a fundamental gap: while RNNs leverage temporal locality, standard Transformer attention does not.
Transformers have nevertheless revolutionized sequential modeling, achieving state-of-the-art results across language and time-series benchmarks. At the core of their success lies the multihead attention (MHA) mechanism, which integrates information from all positions through pairwise similarity. This all-to-all formulation is powerful for natural language processing (NLP), where dependencies between words can span long distances and do not depend strongly on order. Yet the same property can be a liability for time-series forecasting and structured sequential tasks, where local and causal structure is essential, and less relevant distant interactions can obscure the signal.
Recent work has shown that introducing recency biases can improve Transformer performance on weakly local and even noncausal language tasks (Press et al., 2022; Raffel et al., 2023; Liu et al., 2025). These results suggest that models benefit from temporal locality even when the underlying data does not demand it. Surprisingly, however, the role of recency bias in time-series forecasting remains underexplored, despite the fact that time-series data is causal and highly local.
In time-series forecasting, each embedding corresponds to a causal temporal measurement, and the strength of dependencies typically decays with temporal separation. Time-series forecasting problems are inherently causal; the input and target sequences do not overlap. However, attention methods in most state-of-the-art models (Nie et al., 2023; Zhou et al., 2022, 2023, 2021; Wu et al., 2021) do not apply a causal mask. Thus, the temporal embeddings are non-causally updated with other input sequence embeddings from later times. This mismatch between the all-to-all nature of Transformer attention and the local-causal structure of time-series data raises a central question: how well suited is standard attention for time-series forecasting?
We develop the Recency Biased Causal Attention (RBCA) framework to generally study the effects of causal and recency biases on Transformers as they are applied to local and causal datasets. We investigate the flexibility of recency biases to simultaneously perturb the attention correlation structure while sufficiently capturing pairwise correlations unique to each dataset. To do so, we leverage insights from time-series forecasting problems to broadly improve upon Transformer’s fundamental capabilities. Specifically, we note that attention weights are a normalized similarity metric similar to correlations. We are further motivated by the observation that many physical systems exhibit heavy-tailed autocorrelations, e.g., the pairwise correlation strength may decay as a power-law as the time delay grows (Clauset et al., 2009).
We develop power-law recency biases to impose a natural correlation structure on the attention weight distribution and systematically compare their effects across tasks. Our study evaluates recency biases in a principled way, highlighting both their strengths and their nuanced differences across causal and locally structured domains in NLP and time-series forecasting. We begin by testing whether recency biases equip Transformers with the basic read, ignore, and write capabilities characteristic of RNNs. Next, we apply RBCA to a standard encoder-decoder Transformer and assess its performance on causal and locally correlated time-series datasets. Finally, we investigate the effects of RBCA on modern time-series forecasting architectures by developing Powerformer, a PatchTST (Nie et al., 2023) variant with RBCA.
@@
  • We propose Recency-Biased Causal Attention (RBCA), a general framework for introducing recency bias into Transformers. With RBCA, we develop power-law-inspired recency biases that induce a heavy-tailed attention distribution while allowing attention to learn the unique data-dependent correlation structure.
  • We evaluate recency biases on the flip-flop task to test core sequential capabilities such as read, write, and ignore. We find that NLP-derived recency biases degrade performance, while our power-law biases enable Transformers to achieve better accuracy.
  • We assess RBCA in time-series forecasting. First, we apply it to a standard Transformer; and then we introduce Powerformer, a simple encoder-only variant that amplifies recency effects. Our power-law biases consistently improve forecasting accuracy across benchmarks. Notably, Powerformer further strengthens the recency bias despite its regularizing effect, thus highlighting the fundamental importance of recency biases for time-series prediction.
@@ 2 RELATED WORK
The success of Transformers (Vaswani et al., 2017) and LLMs inspired Transformer-based time-series models. For example, Time-GPT (Garza et al., 2024) and One-Fits-All (Zhou et al., 2023), which fine-tunes a pretrained GPT-2 (Radford et al., 2019) with temporal embedding. Many other models aimed to alter the attention mechanism or the input embeddings.
Many early time-series Transformer models altered the attention mechanism to improve alignment with time-series properties. Fourier representations separate fast and slow varying dynamics, simplifying attention representations (Zhou et al., 2022; Woo et al., 2022). AutoFormer (Wu et al., 2021) replaces the similarity metric with autocorrelations. FlowFormer (Wu et al., 2022) achieves linear complexity attention without a recency bias. Informer (Zhou et al., 2021) selects high attention weight values, increasing prediction capacity. iTransformer (Liu et al., 2024) inverts the temporal and embedding dimensions during MHA.
Some works alter the input embedding to enrich the input or better align it with NLP tasks. PatchTST patches the input sequence, outperforming more complicated methods (Nie et al., 2023). Chronos (Ansari et al., 2024) tokenizes the continuous input and trains on large datasets. TOTEM (Talukder et al., 2024) also discretizes the input time-series, but uses VQVAE (van den Oord et al., 2017).
To impose a locality bias, recency biases have been used in both NLP (Liu et al., 2025) and time-series tasks. For NLP, ALiBi (Press et al., 2022) employs smooth exponential decay, while block (Lu et al., 2025) and sliding-window (Beltagy et al., 2020) attentions employ a cutoff scheme. Previous recency-biased time-series models include Reformer (Kitaev et al., 2020), which uses a locality-sensitive hashing mechanism, and ETSformer (Woo et al., 2022), which employs a non-learnable exponential decay in one of their attention mechanisms. Earlier time-series CNN-based works also employed temporal dilations to bias towards early data (Li et al., 2019; van den Oord et al., 2016; Franceschi et al., 2019).
Figure 1: Illustration of Recency Biased Causal Attention (RBCA) and how it fits into the Transformer layer and adapts the attention. Panel (a) shows how RBCA replaces MHA in the standard Transformer block. Panel (b) shows a local-causal mask and subsequent weights, where $\Sigma$ corresponds to the softmax function, red and blue indicate positive and negative values, respectively, and white is labeled when ambiguous. Panel (c) demonstrates the 4 different masking functions we apply to the score (left) and their effects on the weights (right).
@@ 3 METHOD
We propose Recency-Biased Causal Attention (RBCA) as a framework for recency biases. In this section, we review standard Multihead Attention (MHA) and demonstrate how we add causal and recency biases to it. Figures 1 and 2 illustrate RBCA and its effect on attention weights.
3.1 Standard Multihead Attention
Transformers typically leverage MHA to compute new embeddings from weighted sums of input embeddings. Let $\mathbf{X}\in\mathbb{R}^{T\times D}$ be a sequence of $T$ input vectors, each of dimension $D$. For all $H$ attention heads, indexed by $h$, MHA first projects $\mathbf{X}$ into query, key, and value matrices, respectively denoted by $\mathbf{Q}_{h}$, $\mathbf{K}_{h}$, and $\mathbf{V}_{h}$: $\mathbf{Q}_{h}=\mathbf{X}\,\mathbf{W}_{h}^{(\text{Q})}$, $\mathbf{K}_{h}=\mathbf{X}\,\mathbf{W}_{h}^{(\text{K})}$, and $\mathbf{V}_{h}=\mathbf{X}\,\mathbf{W}_{h}^{(\text{V})}$. Each projection matrix $\mathbf{W}_{h}^{(\cdot)}\in\mathbb{R}^{D\times D_{k}}$ is learnable, with head size $D_{k}$. The parameters $\mathbf{Q}_{h}$, $\mathbf{K}_{h}$, and $\mathbf{V}_{h}$ thus each have the shape $T\times D_{k}$.
The unnormalized attention similarity scores measure how each query interacts with all keys
$\mathbf{S}_{h}=\frac{\mathbf{K}_{h}\,\mathbf{Q}_{h}^{\top}}{\sqrt{D_{k}}},$ (1)
where the factor of $1/\sqrt{D_{k}}$ helps stabilize the gradients by normalizing the dot products according to the head dimension. In turn, the normalized attention weights $\mathbf{C}_{h}=\text{Softmax}\bigl(\mathbf{S}_{h}\bigr)$ has each row summing to 1. Finally, each head’s output $\tilde{\mathbf{X}}_{h}=\mathbf{C}_{h}\,\mathbf{V}_{h}$ is computed as a weighted sum of the values, concatenated with the other heads, and transformed by a linear projection
@@ 3.1 Standard Multihead Attention
where $\mathbf{W}^{(\text{A})}\in\mathbb{R}^{(H\,D_{k})\times D}$ is a learnable matrix. The result $\tilde{\mathbf{X}}$ has the same dimensionality as the input $\mathbf{X}$ (i.e., $T\times D$), allowing it to be passed into subsequent Transformer layers or decoded.
Figure 2: A comparison between standard non-causal and nonlocal attention (bottom left) and examples of RBCA with varying masks (top row) applied to the same attention score (top left). In the top row, red is maximized at 0 and minimized (white) at -$\infty$. In the bottom row, red is maximized at 1 and minimized (white) at 0. For the attention score red and blue indicate positive and negative.
Enforcing Causality
The standard causal mask prohibits future embeddings from updating past ones:
$\mathbf{M}^{(\text{C})}_{t,t^{\prime}}=\begin{cases}-\infty&t^{\prime}>t,\\ 0&t^{\prime}\leq t,\end{cases}$ (3)
@@
The attention weights $\mathbf{C}^{(C)}_{t,t^{\prime}}=0$ for all $t^{\prime}>t$, ensuring a causal structure for the time-series data.
3.2 Recency-Biased Causal Attention
The recency bias is a mask $\mathbf{M}^{(\text{L})}$ defined by some nonpositive function $f(\Delta t)\leq 0$ that is dependent on the relative time difference ($\Delta t$) between embeddings:
$\mathbf{M}^{(\text{L})}_{i,j}=\begin{cases}0&j>i,\\ f(t_{i}-t_{j})&j\leq i.\end{cases}$ (4)
RBCA combines the recency bias and causal mask:
$\mathbf{M}^{(\text{C,L})}=\mathbf{M}^{(\text{C})}{}$ $+\mathbf{M}^{(\text{L})},\quad\mathbf{S}_{h}^{(\text{C,L})}=\mathbf{S}_{h}+\mathbf{M}^{(\text{C,L})}$ (5)
$\mathbf{C}_{h}^{(\text{C,L})}$ $=\text{Softmax}\bigl(\mathbf{S}_{h}^{(\text{C,L})}\bigr)$ (6)
$C_{h,i,j}^{(\text{C,L})}$ $\propto\begin{cases}0&j>i,\\ \exp\bigl[f(t_{i}-t_{j})\bigr]&j\leq i.\end{cases}$ (7)